# algorithmic

### v1

@@ 4.1 The Informed RLSVI Algorithm for average-reward MDPs
~$\begin{split}\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)&=\frac{1}{2\sigma^{2}}\sum_{k=1}^{L+t}\sum_{h=0}^{H-1}\left(r_{k}+\max_{b}Q^{\theta}_{t+1}(\check{s}_{k},b)-q^{\theta}_{t}(s_{k},a_{k})\right)^{2}\\ &+\frac{1}{2}({\theta_{0:H}}-{\tilde{\theta}_{0:H}})^{\top}\Sigma_{0}({\theta_{0:H}}-{\tilde{\theta}_{0:H}}).\end{split}$
Now, let us consider an imitation learning setting. Let $\tau_{l}=(s_{k},a_{k},\check{s}_{k})_{h=0}^{H-1}$ be the trajectory of the $l$th episode. Let $\hat{\pi}_{t}(a|s)$ denote the empirical estimate of probability of taking action $a$ in state $s$ at time $h$, i.e., an empirical estimate of the expert’s randomized policy. Let $p(\tau)$ denote the probability of observing the trajectory under the policy $\hat{\pi}$.
Let $\pi_{t}^{\beta,\theta}(\cdot|s)$ denote the parametric representation of the policy used by the expert. And let $p^{\beta,\theta}(\tau)$ denote the probability of observing the trajectory $\tau$ under the policy ${\pi}^{\beta,\theta}$. Then, the loss function corresponding to the KL divergence between $\Pi_{l=1}^{L}p(\tau_{l})$ and $\Pi_{l=1}^{L}p^{\beta,\theta}(\tau_{l})$ is given by
~$\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)$ $=D_{KL}\left(\Pi_{l=1}^{L}p(\tau_{l})||\Pi_{l=1}^{L}p^{\beta,\theta}(\tau_{l})\right)=\int\Pi_{l=1}^{L}p(\tau_{l})\log\frac{\Pi_{l=1}^{L}p(\tau_{l})}{\Pi_{l=1}^{L}p^{\beta}(\tau_{l})}$ (23)
$=\sum_{l=1}^{L}\int p(\tau_{l})\log\frac{p(\tau_{l})}{p^{\beta,\theta}(\tau_{l})},=\sum_{l=1}^{L}\sum_{h=0}^{H-1}\log\frac{\hat{\pi}_{t}(a_{k}|s_{k})}{\pi^{\beta,\theta}_{t}(a_{k}|s_{k})}$ (24)
$=\sum_{l=1}^{L}\sum_{h=0}^{H-1}[\log\hat{\pi}_{t}(a_{k}|s_{k})-\log\pi_{t}^{\beta,\theta}(a_{k}|s_{k})]$ (25)
@@
$=-\sum_{l=1}^{L}\sum_{h=0}^{H-1}\left(\beta q^{\theta}_{t}(s_{k},a_{k})-\log\sum_{b}\exp\left(\beta q^{\theta}_{t}(s_{k},b)\right)\right)+\text{constant}.$ (26)
Remark 4.3.
(i) The loss function ~$\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)$ is the same as the second (action-likelihood) term in (22) while the loss function ~$\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)$ is the same as the first and third terms there (except for perturbation) and minus the $\lambda_{2}\beta$ term that corresponds to the prior over $\beta$. (ii) Note that while we used the more common KL divergence for the imitation learning loss function, use of log loss would yield the same outcome.
Thus, the inf-iRLSVI loss function can be viewed as
~$\tilde{\mathcal{L}}(\beta,\theta)=\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)+\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)+\lambda_{2}\beta,$ (27)
thus establishing that the proposed algorithm may be viewed as bridging Online RL with Imitation Learning. Note that the last term corresponds to the prior over $\beta$. If $\beta$ is known (or uniform), it will not show up in the loss function above.
The above also suggests a possible way to generalize and obtain other online learning algorithms that can bootstrap by use of offline datasets. Namely, at each step, they can optimize a general loss function of the following kind:
~$\tilde{\mathcal{L}}_{\alpha}(\beta,\theta)=\alpha\tilde{\mathcal{L}}_{\text{ORL}}(\theta)+(1-\alpha)\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)+\lambda_{2}\beta,$ (28)
where ~$\tilde{\mathcal{L}}_{\text{ORL}}$ is a loss function for an Online RL algorithm, ~$\tilde{\mathcal{L}}_{\text{IL}}$ is a loss function for some Imitation Learning algorithm, and factor $\alpha\in[0,1]$ provides a way to tune between emphasizing the offline imitation learning and the online reinforcement learning.
@@ 5 Conclusions
We then presented an approximate Bayesian online learning algorithm, inf-iRLSVI that can be seen as naturally bridging online reinforcement learning and imitation learning. In prior work (Osband et al., 2019), such algorithms are well-known to be computationally practical in problems that need deep exploration.

### v2

@@ 4.2 Informd RLSVI Bridges Online RL and Imitation Learning
~$\begin{split}\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)&=\frac{1}{2\sigma^{2}}\sum_{k=1}^{N+t}\sum_{h=0}^{H-1}\left(r_{k,h}+z_{k,h}+\max_{a^{\prime}\in\mathcal{A}}Q^{\theta}_{h+1}(\check{s}_{k,h+1},a^{\prime})-Q^{\theta}_{h}(\check{s}_{k,h},a_{k})\right)^{2}\\ &+\frac{1}{2}({\theta_{0:H}}-{\tilde{\theta}_{0:H}})^{\top}\Sigma_{0}({\theta_{0:H}}-{\tilde{\theta}_{0:H}}).\end{split}$ (27)
where $z_{k,h}\sim\mathcal{N}(0,\sigma^{2})$ are random perturbations. The sum of randomized value iteration loss and randomized prior loss in (26) can be seen as an adaptation of (27) for infinite horizon MDP learning problems.
Let $\hat{\pi}_{t}(a|s)$ denote the empirical distribution of actions at state $s$ with all the online and offline data available at time $t$. The KL-divergence imitation learning loss at state $s\in\mathcal{S}$ can be specified by
$\mathcal{L}_{\text{IL}}(\beta,\theta,s)$ $=D_{KL}\left(\hat{\pi}(s)\|\pi^{\beta}(s;q^{\theta})\right)=\sum_{a}\hat{\pi}(a|s)\log\left(\dfrac{\hat{\pi}(a|s)}{\pi^{\beta}(a|s;q^{\theta})}\right)$ (28)
$=\sum_{a}\hat{\pi}(a|s)\log\left(\dfrac{1}{\pi^{\beta}(a|s;q^{\theta})}\right)+\text{constant}$ (29)
@@
$=-\dfrac{1}{\bar{N}(s)}\sum_{t=0}^{N-1}\sum_{a\in\mathcal{A}}\bm{1}_{\{\bar{s}_{k}=s,\bar{a}_{k}=a\}}\log\left(\pi^{\beta}(a|s;q^{\theta})\right)+\text{constant}$ (30)
$=-\dfrac{1}{\bar{N}(s)}\sum_{t=0}^{N-1}\bm{1}_{\{\bar{s}_{k}=s\}}\left(\beta q^{\theta}_{t}(\bar{s}_{k},\bar{a}_{k})-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\beta q^{\theta}_{t}(\bar{s}_{k},a^{\prime})\right)\right)+\text{constant}.$ (31)
where $\bar{N}(s)$ is the visitation count of state $s$ in the offline data.
Therefore, the (pre-randomization) imitation loss in (25) can be written as a weighted sum of the individual imitation loss defined above
[…]
Therefore, we have shown that the randomized loss function (26) can be viewed as the sum of the RLSVI loss related to online learning, and the randomized imitation loss related to imitation learning. Thus the informed RLSVI algorithm can be viewed as bridging online RL with imitation learning.
@@ 5 Conclusions
We then presented an approximate Bayesian online learning algorithm, inf-iRLSVI, that bridges online learning with imitation learning. In prior work (Osband et al., 2019), such algorithms are well-known to be computationally practical in problems that need deep exploration.